OpenStudy (anonymous):

ordering an ice cream cone from a choice of 31 flavors, 3 types of cones, with or without one of the 4 toppings. Determine the different possibilities. (teach me how to solve questions like this)

OpenStudy (anonymous):

I need help approaching a question like this one. so It would be awesome if you could help in explaining how to solve this.

OpenStudy (anonymous):

what are you trying to find out?

OpenStudy (anonymous):

Its called the multiplication principal. If you take the number of options you have for each individual part and multiply them together, you get the total number of possibilities. http://www.icoachmath.com/math_dictionary/Multiplication_Principle.html With four toppings we get 31*3*4 = 372 Without one of the the four toppings we get 31*3 = 93 With or without we add to get 372 + 93 = 465